quadratic equations-word problem
Chris cuts half a rectangular lawn, 40m by 30m, by mowing strips around the perimeter. Jerri cuts the small rectangle left. How wide a strip does Chris cut so that they share the work equally?

Answer by ankor@dixie-net.com(22740)

Chris cuts half a rectangular lawn, 40m by 30m, by mowing strips around the perimeter. Jerri cuts the small rectangle left. How wide a strip does Chris cut so that they share the work equally?
:
FInd the total area of the lawn
40 * 30 = 1200 sq/m
:
Half of the lawn would be 600 sq/m (obviously)
:
Let the width of strip = x
:
The equation using the dimensions of the small rectangle left:
(40-2x)*(30-2x) = 600
:
FOIL
1200 - 140x + 4x^2 = 600
:
Arrange as a quadratic equation:
4x^2 - 140x + 1200 - 600 = 0
4x^2 - 140x + 600 = 0
:
Simplify, divide by 4. (Makes it easier to factor)
x^2 - 35x + 150 = 0
:
Factors to:
(x-30)(x-5) = 0
Two solutions:
x = +30, does not make sense, ignore this one
and
x = +5 ft is our solution
:
:
Check our solution: find the area of half the lawn
(40-10)*(30-10) = 600
:
Did this help make some sense of this for you?